Find the pH of a buffer made ​​of 40 mL of 0.50M HCOOH mixed with 60 mL of 0.40 M HCOONa.
Ka(HCOOH) = 1.6 * 10^-4

I Need full calculation. All the steps. Thanks you very much
1
Expert's answer
2012-10-31T12:28:11-0400
pH = pKa + lg (Csalt/Cacid)
Vtotal=100 ml
Cn1 * V1 = Cn2 * V2
Cn1 = Cn2 * V2 / V1
Cnsalt = 60 * 0.4 /100 = 0.24 N
Cnacid = 40 * 0.5 / 100 = 0.2 N
pKa = - lg(Ka)
pKa = 3.796
pH = 3.796 + lg(0.24/0.2) = 3.875